OpenStudy (anonymous):

Units: 3 manicures in 2 hours; Unit rate: $9.00; Hours 38: What are the estimated earnings?

OpenStudy (anonymous):

3*9=27 (earnings in 2 hours, right? in 38 hours: earning x= 27*38/2=513